OpenStudy (mikezack123):

Solve the equation 7x-17/8 = 4 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

you need to isolate the variable by undo-ing the operation acting on it. Start by what is furthest away from the variable.

OpenStudy (anonymous):

4 = 32/8 7x -17/8 = 32/8 add 17/8 to both sides 7x = 49/8 divide by 7 on both sides x = 49/56 reduce x = 7/8

OpenStudy (mikezack123):

Thanks to the both of you!